OpenStudy (anonymous):

x^2+y^2+6x-18y+65=0 I need to find the radius ?

OpenStudy (anonymous):

(-g, -f) g= half coefficient of x f= half coefficient of y

OpenStudy (anonymous):

\[x^2+6x+y^2-18y=-65\] \[(x+3)^2+(y-9)^2=-65+3^2+9^2=25\] center is (-3,9) radius is 5

OpenStudy (anonymous):

is the centre r = sqrt(g^2+f^2-c)

OpenStudy (anonymous):

g=3 f=-9 c=65 R= sqrt(9+81-65)=5

OpenStudy (anonymous):

Thank you